# Rethinking Improved Privacy-Utility Trade-off with Pre-existing Knowledge for DP Training

### v1

@@ V-B1 Convergence Analysis
V-B2 Model Accuracy
Table I shows comparative results with prior works. To be fair, we compare the test accuracy of the trained models under the constraint of identical $\epsilon$. We can see that $\mathsf{DP\text{-}Hero}$ improves the test accuracy of state-of-the-arts [14, 52, 53, 54, 55, 16, 56, 46]. In most cases, our $\mathsf{DP\text{-}Hero}$ SGD could attain $>98\%$ test accuracy on the MNIST dataset, whereas other works achieve $95\%\sim 97\%$. Only several works were trained over the CIFAR-10 dataset, yet with the $<60\%$ accuracy. In contrast, $\mathsf{DP\text{-}Hero}$ SGD could achieve $>64.5\%$ accuracy, showing much better results.
Specifically, $\mathsf{DP\text{-}Hero}$ SGD improves $18\%$ accuracy on [55], $47\%$ accuracy on [16], and $22\%$ accuracy on [54]. Training a DP model over the CIFAR-10 dataset may require a pretraining phase, whereas $\mathsf{DP\text{-}Hero}$ SGD training could alleviate this. It shows that $\mathsf{DP\text{-}Hero}$ SGD behaves better on more representative datasets (e.g., CIFAR-10$>$MNIST) than DP-SGD. Figure 4 shows a box-whisker plot on accuracy given varying $\epsilon$. Except for following identical configuration of $\epsilon$, we show additional results with $\epsilon=1,2,3,4$. The test accuracy is relatively stable for different $\epsilon$ in different training processes. When $\epsilon$ is very large, although test accuracy is high, DP protection may not be sufficient for practical usage. Experimental results show that $\mathsf{DP\text{-}Hero}$ SGD is more robust against large noise and supports faster convergence, especially for representative datasets.
TABLE I: Test Accuracy Compared with Prior Top-tier Works
@@ V-B2 Model Accuracy
$2$ $95\%$ $\mathbf{98.10\%}$
$8$ $97\%$ $\mathbf{98.11\%}$
Abadi et al. [14] $\infty$ ${98.3\%}$ $98.12\%$
@@
Bu et al. [53] $5.07$ $97.0\%$ $\mathbf{98.10\%}$
Chen et al. [54] $2.5$ $90.0\%$ $\mathbf{98.12\%}$
Nasr et al. [55] $3.2$ $96.1\%$ $\mathbf{98.11\%}$
Yu et al. [16] $6.78$ $93.2\%$ $\mathbf{98.10\%}$
$1$ $95.82\%$ $\mathbf{98.11\%}$
Ghazi et al. [56] $2$ $98.78\%$ $98.10\%$
@@
MNIST Tramer et al. [46] $2.5,2.9$ $\approx 98\%$ $\mathbf{98.12\%}$
Nasr et al. [55] $3.0$ $55.0\%$ $\mathbf{64.93\%}$
Yu et al. [16] $6.78$ $44.3\%$ $\mathbf{65.04\%}$
CIFAR-10 Chen et al. [54] $8.0$ $53.0\%$ $\mathbf{65.12\%}$
@@ V-C Explaining Experiments
V-C1 Tracking Initial-Phase Training
@@
V-C2 Visualizing DP Training
Given high-resolution and precise class discrimination, we apply Grad-CAM [48] to show visual results on DP training. In Grad-CAM [48], the class-discriminative localization map of width $u$ and height $v$ for any class $c$ is defined to be $\mathcal{L}_{\mathsf{GradCAM}}=\mathsf{ReLU}(\sum_{k}\alpha_{k}^{c}A^{k})$. Here, the weight $\alpha_{k}^{c}$ represents a partial linearization of the downstream feature map activation $A$. In our experiments, we adopt GridCam [48] for interpreting/visualizing how DP noise affects model training. In a model training process, GridCam is employed to visualize explanations of learning features, with or without DP noise.
GridCam [48] can coarsely locate the important regions in the image for predicting the concept, e.g., “dog” in a classification network. Figure 7 visualizes the heat map of training with $\mathsf{DP\text{-}Hero}$ SGD compared with Figure 8. $\mathsf{DP\text{-}Hero}$ SGD training still maintains the representation ability to locate the important objects. That is, the reason for more satisfying accuracy is that the noise added to the gradients could not affect on models’ ability for relatively accurate visualization in a statistical manner, i.e., still protecting individual privacy.
V-C3 A Practical View of Privacy Parameters
@@
$\mathsf{DP\text{-}Hero}$ SGD assigns $\sigma$ as a variable during DP training. As for unbiased noise distribution, $\mu=0$ holds at every execution of sampling noise. In probability theory, the sum of multiple independent normally distributed random variables is normal, and its variance is the sum of the two variances. We use this conclusion to assign the $\sigma$ over diverse features at each training iteration $t$. If we regard all assigned $\sigma$ at each iteration as a matrix, all entries in this matrix vary at different iterations. The parameter configuration at every iteration follows Theorem 3, supporting linearity relation to value in $\mathsf{SVec}$ in $\mathsf{DP\text{-}Hero}$ SGD. Although the theoretical expectation of introducing Gaussian noise with 00 mean value remains identical to the clean model, practical training shifts the expected results to some extent.
@@
V-C4 Understanding of Improved Model Convergence
Motivated by utility improvement, we perform repeated experiments similar to §$\S$ V-B to attain the relation between model training and noise heterogeneity in empiricism. We repeatedly train an identical model given various heterogeneity (adjust noise scales to diverse model parameters for early-stage tests) and witness the corresponding phenomenon in the convergence process. Pure SGD training could attain the best accuracy and converge fastest while training with DP-SGD slows down the convergence constrained by identical remaining configurations. Even after the model’s convergence, DP-SGD training can not reach the highest accuracy as pure SGD training.
@@ V-D Further Discussion
VI Conclusion
@@
Through theoretical and empirical understanding of privacy-utility space, we extend the research line of improving training performance for DP learning by designing a plug-in optimization for training with DP-SGD. The proposed DP-Hero is a versatile differential privacy framework incorporating a heterogeneous DP noise allocation manner. The primary innovation of DP-Hero is its ability to utilize the knowledge embedded in previously trained models to guide the subsequent distribution of noise heterogeneity, thereby optimizing its utility. Building on the foundation of DP-Hero, we introduce a heterogeneous version of DP-SGD, in which the noise introduced into the gradients varies. We have carried out extensive experiments to validate and elucidate the efficacy of the proposed DP-Hero. Accordingly, we provide insights on enhancing the privacy-utility space by learning from the pre-existing leaked knowledge encapsulated in the previously trained models.
We point out a new way of thinking about model-guided noise allocation for optimizing SGD-dominated convergence under the DP guarantee. Besides, we explore explaining DP training via visual representation, reasoning the improved utility. Such an explainable view could benefit from understanding DP protection more vividly, for potentially being against attacks better. In a broader context, we expect heterogeneous DP learning to be adopted beyond (DP-)SGD-based instantiations.

### v3

@@ Introduction
Model Accuracy
Table 1 shows comparative results with prior works. To be fair, we compare the test accuracy of the trained models under the constraint of identical $\epsilon$. We can see that $\mathsf{DP\text{-}Hero}$ improves the test accuracy of state-of-the-arts (Abadi et al. 2016; Feldman and Zrnic 2021; Bu et al. 2019; Chen and Lee 2020; Nasr, Shokri, and Houmansadr 2020; Yu et al. 2019; Ghazi et al. 2021; Tramèr and Boneh 2021). In most cases, our $\mathsf{DP\text{-}Hero}$ SGD could attain $98\%$ test accuracy on the MNIST dataset, whereas other works achieve $95\%\sim 97\%$. Only several works were trained over the CIFAR-10 dataset, yet with the $60\%$ accuracy. In contrast, $\mathsf{DP\text{-}Hero}$ SGD could achieve $64.5\%$ accuracy, showing much better results.
Specifically, $\mathsf{DP\text{-}Hero}$ SGD improves $18\%$ accuracy on (Nasr, Shokri, and Houmansadr 2020), $47\%$ accuracy on (Yu et al. 2019), and $22\%$ accuracy on (Chen and Lee 2020). Training a DP model over the CIFAR-10 dataset may require a pretraining phase, whereas $\mathsf{DP\text{-}Hero}$ SGD training could alleviate this. It shows that $\mathsf{DP\text{-}Hero}$ SGD behaves better on more representative datasets (e.g., CIFAR-10>>MNIST) than DP-SGD. Figure 4 shows a box-whisker plot on accuracy given varying $\epsilon$. Except for following identical configuration of $\epsilon$, we show additional results with $\epsilon=1,2,3,4$. The test accuracy is relatively stable for different $\epsilon$ in different training processes. When $\epsilon$ is very large, although test accuracy is high, DP protection may not be sufficient for practical usage. Experimental results show that $\mathsf{DP\text{-}Hero}$ SGD is more robust against large noise and supports faster convergence, especially for representative datasets.
When extending to FedFed, accuracy is sensitive to the degree of data heterogeneity (non-IID-ness) among clients. Under the IID scenario, where each client’s local data distribution mirrors the global distribution, the global model can closely match the centralized $98\%$ approximately over MNIST, typically reaching $97\%$-$98\%$ accuracy. However, as data becomes more heterogeneous, such as when clients have unbalanced or partially disjoint label distributions, overall accuracy declines. Moderate non-IID partitions (Dirichlet distribution parameter $\alpha>10$) generally yield $95\%$-$97\%$ accuracy, while higher heterogeneity ($\alpha<10$) can reduce accuracy to $90\%$-$95\%$.
Table 1: Test Accuracy Compared with Prior Top-tier Works
@@
22 $95\%$ $\mathbf{98.10\%}$
88 $97\%$ $\mathbf{98.11\%}$
DP-SGD $\infty$ $98.3\%$} $98.12\%$
@@
GDP $5.07$ $97.0\%$ $\mathbf{98.10\%}$
SAS-DP $2.5$ $90.0\%$ $\mathbf{98.12\%}$
DP-GED $3.2$ $96.1\%$ $\mathbf{98.11\%}$
DP-MP $6.78$ $93.2\%$ $\mathbf{98.10\%}$
11 $95.82\%$ $\mathbf{98.11\%}$
LDP 22 $98.78\%$ $98.10\%$
@@
MNIST FeatureDP $2.5,2.9$ $\approx 98\%$ $\mathbf{98.12\%}$
DP-GED $3.0$ $55.0\%$ $\mathbf{64.93\%}$
DP-MP $6.78$ $44.3\%$ $\mathbf{65.04\%}$
CIFAR-10 SAS-DP $8.0$ $53.0\%$ $\mathbf{65.12\%}$
@@
Tracking Initial-Phase Training
@@
Visualizing DP Training
Given high-resolution and precise class discrimination, we apply Grad-CAM (Selvaraju et al. 2017) to show visual results on DP training. In Grad-CAM (Selvaraju et al. 2017), the class-discriminative localization map of width $u$ and height $v$ for any class $c$ is defined to be $\mathcal{L}_{\mathsf{GradCAM}}=\mathsf{ReLU}(\sum_{k}\alpha_{k}^{c}A^{k})$. Here, the weight $\alpha_{k}^{c}$ represents a partial linearization of the downstream feature map activation $A$. In our experiments, we adopt GridCam (Selvaraju et al. 2017) for interpreting/visualizing how DP noise affects model training. In a model training process, GridCam is employed to visualize explanations of learning features, with or without DP noise.
GridCam (Selvaraju et al. 2017) can coarsely locate the important regions in the image for predicting the concept, e.g., “dog” in a classification network. Figure 7 visualizes the heat map of training with $\mathsf{DP\text{-}Hero}$ SGD compared with Figure 8. $\mathsf{DP\text{-}Hero}$ SGD training still maintains the representation ability to locate the important objects. That is, the reason for more satisfying accuracy is that the noise added to the gradients could not affect on models’ ability for relatively accurate visualization in a statistical manner, i.e., still protecting individual privacy.
A Practical View of Privacy Parameters
@@
$\mathsf{DP\text{-}Hero}$ SGD assigns $\sigma$ as a variable during DP training. As for unbiased noise distribution, $\mu=0$ holds at every execution of sampling noise. In probability theory, the sum of multiple independent normally distributed random variables is normal, and its variance is the sum of the two variances. We use this conclusion to assign the $\sigma$ over diverse features at each training iteration $t$. If we regard all assigned $\sigma$ at each iteration as a matrix, all entries in this matrix vary at different iterations. The parameter configuration at every iteration follows Theorem 3, supporting linearity relation to value in $\mathsf{SVec}$ in $\mathsf{DP\text{-}Hero}$ SGD. Although the theoretical expectation of introducing Gaussian noise with 0 mean value remains identical to the clean model, practical training shifts the expected results to some extent.
@@
Understanding of Improved Model Convergence
Motivated by utility improvement, we perform repeated experiments similar to $\S$ Model Utility and Training Performance to attain the relation between model training and noise heterogeneity in empiricism. We repeatedly train an identical model given various heterogeneity (adjust noise scales to diverse model parameters for early-stage tests) and witness the corresponding phenomenon in the convergence process. Pure SGD training could attain the best accuracy and converge fastest while training with DP-SGD slows down the convergence constrained by identical remaining configurations. Even after the model’s convergence, DP-SGD training can not reach the highest accuracy as pure SGD training.
@@
Related Works
Differential Privacy for Deep Learning
Differential privacy has emerged as a solid solution to safeguard privacy in the field of deep learning. Differential privacy (DP) for deep learning can be classified into four directions: input perturbation (Duchi, Jordan, and Wainwright 2013; Shokri and Shmatikov 2015), output perturbation (Papernot et al. 2018), objective perturbation (Chaudhuri, Monteleoni, and Sarwate 2011; Iyengar et al. 2019), and utility optimization (McMahan and Andrew 2018; Abadi et al. 2016; Chen, Wu, and Hong 2020; Yu et al. 2019), showing valuable insights in the aspects of theory and practice. DP could quantify to what extent individual privacy (i.e., whether a data point contributes to the training model) in a statistical dataset is preserved while releasing the established model trained over the specific dataset. Typically, DP learning has been accomplished by applying the unbiased Gaussian noise to the gradient descent updates, a notable example of DP-SGD (Abadi et al. 2016). To be specific, DP-SGD adds the i.i.d. noise sampled from Gaussian distribution to model gradients to protect example-level training data involved in the training process in every iteration.
The noise-adding mechanism has been widely adopted in various learning algorithms, e.g., image classification and natural language processing. PATE (Papernot et al. 2018) is an approach to providing differentially private aggregation of a teacher-student model. Due to the property of post-processing (Dwork and Roth 2014), the student’s model is differentially private since it trains over the noisy inputs. Bayesian differential privacy (Triastcyn and Faltings 2020) takes into account the data distribution for practicality (Jagielski, Ullman, and Oprea 2020). By instantiating hypothetical adversaries (Nasr et al. 2021), various threat models are employed to show corresponding levels of privacy leakage from both the views of practitioners and theoreticians.
Privacy auditions and attacks, or cryptographic protection belong to orthogonal research directions, focusing on the evaluative estimation of the privacy guarantee or cipher-text transmission. Membership inference attack (Ye et al. 2022) enables detecting the presence or absence of an individual example, implying a lower bound on the privacy parameter $\epsilon$ via statistics (Esmaeili et al. 2021). Notably, Steinke et al. (Steinke, Nasr, and Jagielski 2023) theoretically proves the feasibility of auditing privacy through membership inference on multiple examples simultaneously, elaborating an efficient one-round solution. Combining different techniques with this work can be promising, while it is out of scope for this work.
Privacy-Utility Trade-off
For acquiring higher utility (Schäler, Hütter, and Schäler 2023), recent works explore the adaptive mechanism of DP training from different perspectives. They try to either reallocate/optimize the privacy budget (Lee and Kifer 2018; Mohammady et al. 2020; Yu et al. 2019; Geng and Viswanath 2016; Luo et al. 2021) or modify the clip-norms (Pichapati et al. 2019; van der Veen et al. 2018) of a (set of) fixed noise distribution(s) in each iteration. Such a branch of work points out a promising direction of adaptivity via redesigning the randomness. Privacy budget scheduling (Luo et al. 2021) improves the utility of differentially private algorithms in various scenarios. Unlike the aforementioned advances of dynamic noise allocation, our exploration of adjusting noise heterogeneity by model parameters aims to improve the utility of the established model at every iteration rather than optimizing the noise allocation in the range of the whole training process with a constant budget. Handcrafted features, learned from public data, can improve model utility given a fixed privacy budget (Tramèr and Boneh 2021). Rather than introducing auxiliary data, we aim to extract knowledge from protected model parameters without extra data assistance.
Previous analyses have enabled an understanding of utility bounds for DP-SGD mainly in an empirical manner. Altschuler and Talwar (Altschuler and Talwar 2022) explored the theory foundation of privacy loss – how sensitive the output of DP-SGD is. They solve a tighter utility bound given the privacy loss as a function of the number of iterations, concluding that after a small burn-in period, running DP-SGD longer leaks no further privacy. In this work, we exploit visual explanation (Selvaraju et al. 2017) and theoretical understanding to explore the essence of privacy-utility space.
Conclusion
@@
Through theoretical and empirical understanding of privacy-utility space, we extend the research line of improving training performance for DP learning by designing a plug-in optimization for training with DP-SGD. The proposed $\mathsf{DP\text{-}Hero}$ is a versatile differential privacy framework incorporating the heterogeneous DP noise. The primary innovation of $\mathsf{DP\text{-}Hero}$ is its ability to utilize the knowledge embedded in previously trained models to guide the subsequent distribution of noise heterogeneity, thereby optimizing its utility. Building on the foundation of $\mathsf{DP\text{-}Hero}$, we introduce a heterogeneous version of DP-SGD, in which the noise introduced into the gradients varies. We have carried out extensive experiments to validate and elucidate the efficacy of $\mathsf{DP\text{-}Hero}$. Accordingly, we provide insights on enhancing the privacy-utility space by learning from the pre-existing leaked knowledge encapsulated in the previously trained models. Broadly, we point out a new way of thinking about model-guided noise allocation for optimizing SGD-dominated convergence under the DP guarantee. Besides, we explore explaining DP training via visual representation, reasoning the improved utility. Such an explainable view could benefit from understanding DP protection more vividly, for potentially being against attacks better.